OpenStudy (anonymous):

Anyone here taking Geometry A in Connect Academy/Connexus and is past Unit 4; Lesson 9?

jigglypuff314 (jigglypuff314):

Hello most prospective tutors and I would be more inclined to answer you quickly if you'd please post your actual question. then anyone willing to help you will help you. :) if, on the other hand, you did this in order to find someone to cheat off of, please know that OpenStudy has an Anti-Cheating policy and cheating is not allowed

OpenStudy (anonymous):

Everytime I post questions on here no one wants to help me no matter how much I beg them. They're too busy chatting.

jigglypuff314 (jigglypuff314):

I can try to help, post your question and I'll take a look at it :)

OpenStudy (anonymous):

Given ^RST = ^NPQ, m<R = -7x + 9, m<N = -10x, find m<R and M<N? 30 60 40 50

jigglypuff314 (jigglypuff314):

RST = NPQ so angle R = angle N so -7x + 9 = -10x can you solve for x from there? :)

OpenStudy (anonymous):

No, that's why I need help. I stink at Geometry so much that I'm failing for the 4th time.

jigglypuff314 (jigglypuff314):

-7x + 9 = -10x first add 10x to both sides +10x +10x and get 3x + 9 = 0 then subtract 9 from both sides -9 -9 3x = -9 then divide both sides by 3 /3 /3 x = -9/3 = ? do you see how I got all that?

OpenStudy (anonymous):

Yeah, I just somehow get it wrong everytime I try lol. Ok so x = -9/3 = -3?

jigglypuff314 (jigglypuff314):

yep ^_^

jigglypuff314 (jigglypuff314):

then plug that back into m<R = -7x + 9 then m<N = -10x

OpenStudy (anonymous):

30

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

Ok, can you please stay and don't leave me lol. I have more questions I need help in.

jigglypuff314 (jigglypuff314):

sure ^_^

OpenStudy (anonymous):

Which of the following statements is true about the triangles below(give me a minute to draw that). ^ABC = ^DEF by SSA ^ABC = ^DEF by AAS ^ABC = ^DEF by SAS ^ABC = ^DEF by ASA

OpenStudy (anonymous):



jigglypuff314 (jigglypuff314):